Determining Amounts to be Paid on Invoices

Determine the amount to be paid in full settlement of each of the following invoices, assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period:

Merchandise
Invoice AmountFreight
Paid by Seller Customer Returns
and Allowances

a.$14,400           —             FOB destination, n/30$700             

b.12,000           $500            FOB shipping point, 2/10, n/301,400             

c.5,800           —             FOB shipping point, 1/10, n/30500             

d.4,400           200             FOB shipping point, 1/10, n/30600             

e.3,500           —             FOB destination, 2/10, n/30—             

a. $

b. $

c. $

d. $

e. $
Answer #1
Amount paid = Invoice Amount+Freight Paid -Returns and Allowances-Discount
a 13700 =14400-700
b 10888 =(12000-1400)*(1-2%)+500
c 5247 =(5800-500)*(1-1%)
d 3962 =(4400-600)*(1-1%)+200
e 3430 =3500*(1-2%)